OpenStudy (anonymous):

Which is a solution of the inequality? p + 4 – 2(p – 10) > 0 A. 29 B. 26 C. 24 D. 22

OpenStudy (anonymous):

D 22 If you're not sure on a problem like this, plug in the answers using order of operations, so for answer choice D, start with the parenthesis: 2(p-10) Plug in 22 for P and solve: (22-10) = 12 12*2=24 So for this answer choice, you have solved the part of the equation that says: 2(p-10), it is 24. Now plug in 22 for the other variable and solve the equation. 22+4-24= 2 2>0

OpenStudy (anonymous):

I realize that I solved that inequality as an equation - that's fine for this case because you're only looking for an answer that is a positive number (greater than 0)